what is the total electric flux due to these two point charges through a spherical surface centered at the origin and with radius r 1 = 0.610 m ? Constants A point charge q.-3.95 nC is located on the x-axis at z 2.25 m, and a second point charge g2--5.50 C is on the y-axis at y 1.25 m Submit Request Answer ▼ Part B What is the total electric flux due to these two point charges through a spherical surface centered at the origin and with radius r2 1.40 m? Φ= Submit Part C What is the total electric flux due to these two point charges through a spherical surface centered at the origin and with radius a 2.40 m? 重- Submit t An
Answer #1

Solution::

φ = Q /ξ0
a) No charge is enclosed. φ = 0

b) Charge enclosed = -5.50 nC
φ = Q /ξ0 = -5.50 x 10^-9 / 8.854 x 10^-12
= - 621.188 Cm/F
Negative sign indicate flux enters the surface.

c) Charge enclosed = [3.95 -5.50] nC

φ = Q /ξ0 = - 1.55 x 10^-9 / 8.854 x 10^-12
= -175.06 Cm/F